# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v2

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly being used for odometry and SLAM applications. However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing point cloud matching based solutions, especially those originally intended for more accurate sensors such as LiDAR. Inspired by visual odometry research around 3D Gaussian Splatting [10], in this paper we propose using freely positioned 3D Gaussians to create a summarized representation of a radar point cloud tolerant to sensor noise, and subsequently leverage its inherent probability distribution function for registration (similar to NDT [2]). Moreover, we propose simultaneously optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function. Finally, we fuse our Gaussian modeling and scan matching algorithms into an EKF radar-inertial odometry system designed after current best practices. Experiments using publicly available 4D radar datasets show that our Gaussian-based odometry is comparable to existing registration algorithms, outperforming them in several sequences.
@@ I Introduction
4D millimeter-wave (mmWave) radars have earned considerable popularity in the world of robotics as an alternative to classic camera or LiDAR sensors due to their robustness against adverse weather conditions (rain, snow, fog, etc.), as well as their compact size and low power consumption. As such, there is considerable interest in incorporating these sensors in odometry and SLAM applications. Existing LiDAR odometry solutions have been applied, adapted and sometimes even further specialized to 4D radar, with varying degrees of success [28, 29, 30, 26, 24]. However, the point clouds generated by 4D radars are sparser, noisier, and have a narrower field of view compared to those generated by LiDARs; and as a result these algorithms often struggle with the geometry of certain scenes. Therefore, it is desirable to explore scene representations that directly model coarse-grained surfaces and regions of radar point clouds instead of focusing on individual points themselves.
Another technique that has gained popularity recently in the field of robotics is 3D Gaussian Splatting (3DGS) [10]. This algorithm creates an explicit representation of a visual scene, and has been applied to SLAM [16, 9, 1] as an alternative to existing methods using implicit representations such as NeRF [17]. However, 3DGS heavily relies on the nature of visual information for training and rendering, and thus adapting it for use with LiDAR or radar instead of visual cameras is challenging; as opposed to NeRF, which can be straightforwardly adapted to model a SDF instead [18].
@@
In this paper we draw inspiration from both lines of research to propose the use of Gaussian models for representing a point cloud and performing registration with respect to it, as an alternative to existing registration methods such as the Normal Distributions Transform (NDT) [2] or Generalized Iterative Closest Point (GICP) [21]. Gaussians are freely positioned in space instead of following a regular grid, as in 3DGS and unlike NDT. Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can in some situations increase the performance of the system, as we show in our experiments. In addition, we outline an Extended Kalman Filter (EKF) based approach to radar-inertial odometry that consolidates existing practices by integrating both radar egovelocity [4] and keyframing/scan matching components, which we then use to evaluate our Gaussian-based approach in comparison to other registration algorithms on multiple datasets and radar types.
Figure 2: Overview of the proposed Gaussian radar-inertial odometry pipeline
Figure 2 shows the overall pipeline of our odometry system. As in other solutions, we use an Extended Kalman Filter (EKF) to perform inertial strapdown and process sensory input, including radar egovelocity. Keyframing is used to select individual 4D radar scans, which then undergo Gaussian modeling. The Gaussian multi-hypothesis scan matching algorithm registers every incoming 4D radar scan against the Gaussian model in order to produce a relative pose observation with respect to the last keyframe, which is then used to further update the EKF.
This paper is organized as follows: Section II reviews existing related literature, Sections III, IV and V describe the three main components of our method respectively: Gaussian modeling, Gaussian scan matching, and the Extended Kalman Filter tying everything together; Section VI shows experimental results, and Section VII closes the paper with some conclusions.
@@ II Related work
II-A Odometry and SLAM
There exist a large number of works focused on odometry, sometimes even attaining a full SLAM pipeline (including loop closure). Particularly we are interested in approaches that use non-visual sensors, such as [27, 25, 19, 4, 5, 28, 29, 30, 26, 24, 31]. These works use a 3D LiDAR or 4D radar sensor, sometimes together with an IMU, to provide an odometry solution. The main technical advantage of these sensors over visual sensors (cameras) is their increased field/depth of view and robustness to adverse environmental conditions: in particular, both are immune to low light conditions, and 4D radar in addition offers robustness to fog, dust and rain. Moreover, 4D radar is uniquely capable of measuring the velocity of objects by leveraging the Doppler effect. Researchers have taken advantage of this to directly measure the robot’s own velocity (“egovelocity”) [4, 5], something which has been shown to significantly improve the performance of odometry systems.
@@ II-A Odometry and SLAM
Registration of incoming point clouds with respect to previous ones is a key technique used in some of the cited odometry systems. In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [21]. This algorithm calculates distribution probability at each point based on their neighborhood, and optimizes a distribution-to-distribution scan matching function. In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position. For this reason [30] proposed the Adaptive Probability Distribution variant of the algorithm (APDGICP), which leverages GICP’s ability to make use of the spatial probability distribution of each point, calculating the required point position covariance matrices according to the radar specifications. However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
The works cited in this section use different sources of information, however many do not leverage all mentioned sources at the same time. Table I includes a list of surveyed methods.
II-B Gaussian-based modeling
@@
The Normal Distributions Transform (NDT) algorithm [2] slices and subdivides a point cloud using a regular grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan. An important limitation of this algorithm is the fact that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in the fact that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. Moreover, the original authors only consider 2D scan matching (requiring 3D point clouds to be projected into the XY plane), although subsequent works [15] extend the algorithm to 3D. NDT has been used with radar scans too [13, 14], albeit without inertial information. Another similar algorithm is the Voxelized GICP (VGICP) [12], which also creates an NDT-like voxel-based representation to serve as the reference model, and combines it with the distribution-to-distribution based approach of GICP.
@@
3D Gaussian Splatting (3DGS) [10] is a popular technique for creating a model of a scene from multiple posed images. A 3DGS model is a collection of multiple 3D normal distributions (Gaussians) freely distributed in space, each also containing color/radiance information encoded using spherical harmonics. Even though this technique was originally developed as an alternative to NeRF [17] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [16, 9, 1]. However, 3DGS and these works are all designed for visual sensors (cameras), and they cannot leverage other sensors such as LiDAR or radar due to the nature of the representation. This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [18] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference; thus requiring significant redesign efforts in order to synthesize other types of data such as LiDAR scans [3].
@@ III Gaussian modeling
Given a point cloud $\mathcal{P}$ contaning $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization. Following the line of works such as NDT [2] and 3DGS [10], we decided to model the geometry using normal distributions. Figure 1 shows an example of a Gaussian model.
@@ III-C Optimization
We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF of the multivariate normal distribution:
@@ IV-C Keyframing
A common technique used in odometry systems is keyframing. Certain poses of the robot are selected to become local reference points (keyframes) that are later used for registration, graph optimization or loop closure purposes. In our case, we use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds. This has the benefit of reducing the accumulated error that arises from successive applications of the scan matching algorithm, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate the appropriate keyframe pacing. We use a simple set of criteria to decide when a new pose $\xi^{j}$ qualifies to become a new keyframe. Given the most recent keyframe $\xi^{i}$, we can derive the $SE(3)$ transformation relative to the keyframe $T=\xi^{j}\ominus\xi^{i}=\{\mathbf{t},\mathbf{q}\}$. The criteria are then simply that any of the following must hold:
  • For the translational component: $||\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
@@ V Inertial odometry with EKF
We use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results. Unlike other published solutions [22, 4, 5], we only use Error State filtering for rotational variables, as opposed to the entire state. We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$). The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$. We decompose each $\mathbf{C}$ into an error part and a nominal quaternion part as follows: $\mathbf{C}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\mathbf{R}\{\mathbf{q}\}$, approximating the error part as $\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization.
@@
[…]
@@ V-E Scan matching observation
Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching
@@
(Ablation) Our EKF with NDT [2] 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
(Ablation) Our EKF with GICP [21] 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
(Ablation) Our EKF with VGICP [12] 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
(Ablation) Gaussian single hypothesis 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
Gaussian multi-hypothesis 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
@@
(Ablation) Our EKF with NDT [2] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
(Ablation) Our EKF with GICP [21] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
(Ablation) Our EKF with VGICP [12] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
(Ablation) Gaussian single hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
@@
(Ablation) Our EKF with GICP [21] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
(Ablation) Our EKF with VGICP [12] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
(Ablation) Gaussian single hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
@@ VI Experimental results
In order to evaluate the performance of our Gaussian modeling and scan matching method, we use existing public 4D radar datasets. In particular, we selected the NTU4DRadLM [29] and Snail-Radar [8] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [7], while preserving existing practice by using a custom evaluation script that follows the behavior of older packages such as rpg_trajectory_evaluation as closely as possible. In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m because the errors are much smaller.
Besides our full method, we evaluate four ablations within our EKF framework: one without multi-hypothesis registration, and three replacing our Gaussian-based algorithms with NDT, GICP and VGICP. Since these are all single-hypothesis methods, we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm). The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [11] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. In both cases we sample a swarm of 8 particles during scan matching.
The source code of our odometry solution can be accessed online on GitHub222 https://github.com/robotics-upo/gaussian-rio-cpp . We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo. The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, etc.

### v3

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly used for odometry and SLAM (Simultaneous Location and Mapping). However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR. Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent probability distribution function for registration. Moreover, we propose tackling the problem of radar noise by optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function. Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system. Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
@@ I Introduction
4D millimeter-wave (mmWave) radars have earned considerable popularity in the world of robotics as an alternative to classic camera or LiDAR sensors due to their robustness against adverse weather conditions (rain, snow, fog, etc.), as well as their compact size and low power consumption. As such, there is considerable interest in incorporating these sensors in odometry and SLAM applications. A key component of such system is the modeling and scan matching algorithm. Point clouds are used to capture the geometry of the scene at a given time, and subsequent scans are compared (matched) to the reference model in order to extract localization information. Popular established algorithms that perform this task include the Normal Distributions Transform (NDT) [1] or Generalized Iterative Closest Point (GICP) [2]. However, despite working well for LiDAR [3, 4, 5], radar point clouds are considerably more challenging to register because they are sparser, noisier, and have a narrower field of view; and, as a result, these algorithms often struggle with the geometry of certain scenes.
@@
Figure 2: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes
We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians. Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene (see Fig. 1). As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians. This approach is entirely featureless, and does not require any priors other than an initial set of centroids. Moreover, we propose a registration step that considers multiple simultaneous hypotheses in order to add further robustness against radar noise. Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig. 2), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V describes the EKF-based RIO pipeline used in the experiments; Section VI shows experimental results, and Section VII concludes the paper proposing future work.
@@ II Related work
Registration of incoming point clouds with respect to previous ones is a key technique used in 3D LiDAR and 4D radar odometry and SLAM systems [3, 4, 5, 7, 8, 9, 10, 11, 12]. In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [2]. This algorithm calculates a distribution probability at each point based on its neighborhood, and optimizes a distribution-to-distribution scan matching function. In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position. For this reason [9] proposed the Adaptive Probability Distribution variant of the algorithm (APDGICP), which leverages GICP’s ability to make use of the spatial probability distribution of each point, calculating the required point position covariance matrices according to the radar specifications. However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
@@
The Normal Distributions Transform (NDT) algorithm [1] subdivides a point cloud using a voxel grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan. An important limitation of this algorithm is that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. NDT has been used with radar scans too [13, 14]. Voxelized GICP (VGICP) [15] is a derivative that combines NDT’s voxel-based representation with the distribution-to-distribution based approach of GICP for registration. G3Reg [16] is another derivative of NDT, which performs classification on each voxel (plane/cluster/point) in order to merge plane cells together, and extract semantic information to further aid scan matching. In all these algorithms, each individual Gaussian is modeled separately using their associated subcloud extracted from the segmentation. However, we are interested in approaches that jointly optimize all Gaussians, looking at the entire point cloud.
@@
3D Gaussian Splatting (3DGS) [6] is a popular technique for creating a model of a scene from multiple posed images. A 3DGS model is a collection of multiple 3D normal distributions (Gaussians) freely distributed in space, each also containing color/radiance information encoded using spherical harmonics. Even though this technique was originally developed as an alternative to Neural Radiance Fields (NeRF) [17] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [18, 19, 20]. Visual-LiDAR methods that improve the robustness of 3DGS with priors sourced from LiDAR input have also been developed [21, 22, 23]. However, 3DGS and these works are all designed for visual input, and they cannot be easily reworked to fully replace it with other sensors such as LiDAR or radar due to the nature of the representation. This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a Signed Distance Field (SDF) [24] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference; thus requiring significant redesign efforts in order to synthesize other types of data such as LiDAR scans [25].
Following 3DGS as inspiration, we define a purely geometric Gaussian modeling algorithm for radar point clouds based on simultaneous optimization of 3D Gaussians. This representation fundamentally differs from NDT in the fact that Gaussians are not bound to a predefined voxel grid, and all parameters of the model are jointly optimized instead of subdividing the clouds and individually generating a Gaussian for each subcloud.
@@ III Gaussian modeling
Given a point cloud $\mathcal{P}$ contaning $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization. We model the geometry of a given radar scene using normal distributions, as in [1, 6, 15, 16]. Figure 1 shows an example of a Gaussian model.
@@ III-C Optimization
We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF (Probability Density Function) of the multivariate normal distribution:
@@ IV-C Keyframing
A common technique used in odometry systems is keyframing. Certain poses of the robot are selected to become local reference points (keyframes) that are typically used for registration, pose graph optimization or (in the case of SLAM) loop closure purposes. In our case, we only use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds. The system only retains the pose and Gaussian model of the most recently generated keyframe. This approach, compared to trivial successive applications of the scan matching algorithm, has the benefit of reducing the accumulated error, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate an appropriate keyframe pacing. We use a simple set of criteria to decide when a new pose $\xi^{j}$ qualifies to become a new keyframe. Given the most recent keyframe $\xi^{i}$, we can derive the $SE(3)$ transformation relative to the keyframe $T=\xi^{j}\ominus\xi^{i}=\{\mathbf{t},\mathbf{q}\}$. The criteria are then simply that any of the following must hold:
  • For the translational component: $\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
@@ V Inertial odometry with EKF
In order to evaluate the suitability and effectiveness of the Gaussian algorithms for Radar-Inertial Odometry, it is necessary to define a pipeline for use in the experiments. Following well established practice [28, 29, 12], we use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results. In order to simplify the formulation, we only use Error State filtering for rotational variables, as opposed to the entire state. We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$). The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$. We decompose each $\mathbf{C}$ into an error part and a nominal quaternion part as follows: $\mathbf{C}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\mathbf{R}\{\mathbf{q}\}$, approximating the error part as $\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization.
@@
There also exists a vector $\mathbf{g}=-[0\;0\;9.80511]^{T}$ modeling the Earth’s gravity treated as a constant, as we do not consider long enough travel distances for changes in the plane tangent to the Earth’s surface to matter.
@@ V-E Scan matching observation
Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching, Abl.: Ablation
@@
(Abl.) NDT Scan Matching [1] 12.61 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
(Abl.) GICP Scan Matching [2] 12.33 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
(Abl.) VGICP Scan Matching [15] 6.79 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
(Abl.) Gaussian Single Hypothesis 5.29 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
Gaussian Multi-Hypothesis 19.17 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
@@
(Abl.) NDT Scan Matching [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
(Abl.) GICP Scan Matching [2] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
(Abl.) VGICP Scan Matching [15] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
(Abl.) Gaussian Single Hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
@@
(Abl.) GICP Scan Matching [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
(Abl.) VGICP Scan Matching [15] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
(Abl.) Gaussian Single Hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
@@ VI Experimental results
We use existing public 4D radar datasets for evaluation. In particular, we selected the NTU4DRadLM [8] and Snail-Radar [31] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [32], while preserving existing practice by using a custom evaluation script that follows the behavior of older packages such as rpg_trajectory_evaluation as closely as possible. In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m due to much smaller errors.
Besides our Gaussian methods, we evaluate four ablations within our RIO framework: one without multi-hypothesis registration, and three using common established scan matching algorithms (NDT, GICP and VGICP) directly with radar point clouds. We use the current nominal state of the EKF directly as the initial estimate for single-hypothesis scan matching. The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [33] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. In both cases we sample a swarm of 8 particles during scan matching.
The source code of our system, which includes the Gaussian algorithms and the odometry pipeline, can be accessed online on GitHub [^2] https://github.com/robotics-upo/gaussian-rio-cpp . We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo. The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, etc.